Provide the missing data in the following table for a distributor of martial arts products: (Enter "ROI" answers to one decimal place.)

Division
Alpha Bravo Charlie
Sales ??? $330,000 ?????
Net operating income ???? $39,600 $62,040
Average operating assets $394,000 ???? ?????
Margin 6 % ???? % 11 %
Turnover 5.0 times ???? times ????? times
Return on investment (ROI) ???? % 30.0 % 22.0 %
Answer #1

Margin=Net operating income/Sales

Turnover=Sales/Average operating assets

ROI=Net operating income/Average operating assets

Alpha Bravo Charlie
Sales (394000*5)=$1,970,000 330,000 (62040/0.11)=$564,000
Net operating income (1,970,000*6%)=$118200 39600 62040
Average operating assets 394000 (39600/0.3)=$132,000 (62040/0.22)=$282,000
Margin 6% (39600/330,000)=12% 11%
Turnover 5 (330,000/132000)=2.5 (564000/282000)=2
ROI (118200/394000)=30% 30% 22%